Assume the following endowments of capital and labor for France and Germany: France Germany Capital 50...

  1. Assume the following endowments of capital and labor for France and Germany:

France

Germany

Capital

50

20

Labor

150

100

Both countries have the option to produce either computers or textiles. Computers are capital intensive, while textiles are labor-intensive. Which country will produce computers, and which will produce textiles? Why?
Answer #1

labor capital ratio in France= 150/50= 3

Labor capital ratio in Germany = 100/20= 5

Since labor capital ratio is higher in Germany, Germany should produce labor intensive good.

Thus Germany should produce Textiles and France should produce computers.